# Multimodal Outer Arithmetic Block Dual Fusion of Whole Slide Images and Omics Data for Precision Oncology

### v1

@@ 3.2.2 Ablation studies
For the subtyping task, the core focus of our work is the dual fusion approach. We conducted ablation studies using only MOAD-Net within the Dual Fusion framework, while all other results detailed in Table 1 that are without MOAD-Net indicators showed late fusion outcomes. Moreover, for a comprehensive evaluation of MOAD-Net, we conducted three ablation experiments (detailed in Table 2); In the first experiment, we removed the MOAB fusion block, using only ABMIL. Here, $\mathbf{p}_{ij}$ served as the input to ABMIL, and the resulting feature embedding $\mathbf{W}_{i}$ was directly fed to the classifier layer to assess the distinct impact of both ABMIL and the DNA modality. In the second experiment, we removed the early fusion block (illustrated in Fig. 1), making $\mathbf{e}_{ij}$ the input to the gated attention block, which is then followed by the MOAD-Net fusion block. The results indicate a reasonable overall accuracy, reflecting the complexity of subtype classification; however, they also underscore the need for advanced fusion techniques to leverage complementary features across modalities effectively. Last, we employed a task-agnostic encoder, ConvNext.v2 [35], pre-trained on ImageNet, to extract features from the WSI and tested MOAD-Net with this setup. We also explored an alternative configuration, placing the MOAB fusion block in the early fusion stage, which revealed two primary issues. First, element-wise multiplication in each outer arithmetic operation required reducing both feature dimensions to less than 64 to manage the high patch count (up to 10K per WSI), resulting in substantial information loss and suboptimal performance. Second, this approach demanded computational power and memory resources that exceeded feasible limits due to the large number of patches.
For the survival prediction task, we extensively evaluated the MOAD-Net approach using the two most common baseline fusion models: ABMIL [33] and TransMIL [31]. We adopted two late fusion techniques: concatenation (Cat) and Kronecker product (KP). These were compared with MOAD-Net within both fusion settings, as shown in Table 4.
@@ 3.3 Data and Code Availability
For brain tumor subtyping, we obtained data from BRAIN UK through a rigorous application process, securing anonymized tissue samples and epigenetic data from NHS Neuropathology Centers. For survival prediction, WSIs are publicly available through the TCGA repository, with corresponding omic data accessible from  [^3] https://github.com/mahmoodlab/SurvPath/tree/main/datasets_csv/metadata. Source code will be made available upon acceptance.
@@ 4 Results
4.1 Subtyping and survival prediction
Quantitative subtyping and survival prediction results are depicted in Tables 1 and 3. Overall, MOAD-Net integrated with ABMIL consistently demonstrates superior performance in brain tumor subtyping and BLCA survival prediction, and achieves competitive performance with the best models for BRCA across recent methodologies[23, 19, 13, 20]. For the subtyping task, Table 1 shows that the SNN model performs well across metrics when using omics-only data, closely followed by the MLP model with slightly lower scores. These results indicate that omics data alone offers strong predictive power, though further improvements and interpretability may be limited without incorporating WSI data. Conversely, WSI only models exhibit relatively low performance across metrics, likely due to indistinguishable patterns among most subtypes and the presence of rare classes represented by only a few large slides, making it challenging to obtain sufficient information for high accuracy. These limitations emphasize the importance of multimodal fusion for enhancing interpretability and performance. In a multimodal setting, ABMIL-MOAD-Net achieves the best scores across all metrics, establishing it as the top performing model in the subtyping task and demonstrating the MOAD-Net architecture’s effectiveness in leveraging multimodal data to provide a richer, more informative representation. Notably, integrating MOAD-Net with advanced architectures like TransMIL and ACMIL significantly improves performance (F1-Micro and Recall) compared to using these models without MOAD-Net. Qualitative results in Fig. 4 further illustrate MOAD-Net’s effectiveness in classifying most tumor subtypes, showcasing the strong impact of its intermingled features. To assess class separation accuracy in the t-SNE representations, we calculated silhouette scores for early and late fusion t-SNE plots, as they displayed visually similar patterns. The late fusion t-SNE achieved a silhouette score of 0.33, while MOAD-Net scored 0.37, indicating that MOAD-Net provides more accurate class separation. This is also evident in the box plot in Fig. 3, where MOAD-Net displays fewer low outliers and a more compact distribution than early and late fusion.
Figure 3: Comparison of F1-scores across different fusion methods for glioma classification. The radar chart illustrates the F1-score performance across all 20 classes, highlighting distinct patterns for Early Fusion, MOAD-Net, and Late Fusion. The bar plots zoom in on specific glioblastoma and glioma classes, showing class-level performance variations across fusion methods. The box plot provides a summary of F1-score distributions, showcasing the variability and consistency of each fusion method.
The confusion matrices demonstrate that the early fusion method significantly boosted MOAD-Net’s accuracy in classifying subtypes, especially for minority classes. Notably, classes 6, 9, and 13, which went unclassified with the late fusion method, were correctly identified with early fusion and MOAD-Net, with our method achieving the best performance overall. Conversely, late fusion performed better in classes 2 and 16, highlighting the challenges of handling heterogeneous multimodal data that can lead to misclassification in MOAD-Net. This also underscores the need for tailored fusion techniques to address different data complexities. The clear diagonal in the confusion matrix across all fusion methods confirms the effectiveness of using both early and late fusion techniques within MOAD-Net’s dual fusion framework, emphasizing the distinct contributions of each fusion approach in various scenarios. Furthermore, Fig. 3 illustrates that MOAD-Net’s dual fusion approach significantly enhanced classification of all glioblastoma subtypes (4 groups) and diffuse glioma subtypes (4 groups), even in cases with highly overlapping morphological features.
@@ 4.1 Subtyping and survival prediction
Table 1: Subtyping prediction MOAD-Net results on the UK BRAIN dataset. We show MOAD-Net combined with SOTA baseline models. The best performance is highlighted in bold. Concatenation and Kronecker products are denoted by (Cat) and (KP). The darker gray rows represent models enhanced by integrating MOAD-Net or using its early fusion embeddings $\mathbf{p}_{ij}$, demonstrating improved performance compared to standalone baseline models.
Model F1-Ma F1-Mi Prec Recall
@@
ACMIL MOAD-Net $0.501$$\pm 0.011$ $0.729$$\pm 0.031$ $0.602$$\pm 0.032$ $0.490$$\pm 0.014$
TransMIL (Cat) $0.711$$\pm 0.005$ $0.799$$\pm 0.005$ $0.744$$\pm 0.003$ $0.707$ ±$\pm 0.002$
TransMIL (KP) $0.483$$\pm 0.003$ $0.741$$\pm 0.016$ $0.485$$\pm 0.007$ $0.506$$\pm 0.008$
TransMIL MOAD-Net $0.724$$\pm 0.005$ $0.816$$\pm 0.002$ $0.759$$\pm 0.009$ $0.719$$\pm 0.001$
MCAT $0.402$$\pm 0.046$ $0.661$$\pm 0.021$ $0.408$$\pm 0.045$ $0.414$$\pm 0.039$
MCAT $p_{ij}$ $0.432$$\pm 0.015$ $0.703$$\pm 0.017$ $0.441$$\pm 0.031$ $0.446$$\pm 0.011$
SURVPATH $0.424$$\pm 0.008$ $0.697$$\pm 0.012$ $0.478$$\pm 0.038$ $0.425$$\pm 0.013$
SURVPATH $p_{ij}$ $0.531$$\pm 0.008$ $0.761$$\pm 0.008$ $0.632$$\pm 0.005$ $0.520$$\pm 0.007$
ABMIL (Cat) $0.718$$\pm 0.013$ $0.806$$\pm 0.002$ $0.764$$\pm 0.018$ $0.717$$\pm 0.023$
ABMIL (KP) $0.447$$\pm 0.024$ $0.737$$\pm 0.020$ $0.481$$\pm 0.043$ $0.464$$\pm 0.032$
ABMIL MOAD-Net $\bf 0.745$$\pm 0.025$ $\bf 0.820$$\pm 0.013$ $\bf 0.769$$\pm 0.016$ $\bf 0.745$$\pm 0.035$
@@
Table 2: Subtyping prediction Ablation study of MOAD-Net showing the performance of early, late, and dual fusion methods.
Ablation Model Description F1-Ma
ABMIL Early fusion with $\mathbf{p}_{ij}$ $0.644\pm 0.010$
($-0.10$)
ABMIL - MOAB Late fusion with $\mathbf{e}_{ij}$ $0.690\pm 0.030$
($-0.05$)
MOAD-Net ConvNeXt encoder $0.732\pm 0.012$
($-0.01$)
MOAD-Net Full model $0.745\pm 0.035$
For the survival prediction task, integrating MOAD-Net with existing (SOTA) models (indicated by *) demonstrates a constant performance improvement over their standalone counterparts, emphasizing the value of MOAD-Net’s rich intermingled features. These features enhance the c-index scores across multiple methods, making MOAD-Net particularly effective in predicting patient disease-specific survival for BLCA, where it outperforms all other methods and achieves competitive results in BRCA. Prototype-based methods, such as PIBD and MMP, show notable performance gains on BRCA data compared to other approaches. To further explore this, we integrated MOAB within PIBD [20], achieving results comparable to the leading model [22]. However, we did not test MOAB with MMP [27] because MMP already incorporates two early fusion stages (transformer and optimal transport). Adding MOAB would bring the total to four fusion stages, potentially introducing additional complexity and noise without clear performance benefits. While our method demonstrates promising performance across multiple tasks, it’s essential to acknowledge some limitations. MOAD-Net faces three primary challenges: first, with dual fusion stages, controlling redundant features from omic data becomes limited. These redundancies can potentially hinder the model’s ability to accurately predict certain subtypes, as seen in the confusion matrix where classes 2 and 3 in MOAD-Net score lower than the ablated models shown in Fig. 4B and C. Second, balancing early and late fusion presents a trade-off; although MOAD-Net leverages both stages to capture the strengths of each, finding the optimal balance can be complex. Some cases exhibit performance variations depending on the fusion stage emphasis, complicating model configuration and interoperability. Last, our late fusion block, incorporating MOAB, operates on latent space vectors derived from the early fusion stage, where omics features have already been blended. Consequently, pinpointing the specific CpG feature that most significantly impacts the outcome becomes challenging.
@@
Table 3: Survival prediction Results of MOAD-Net with baselines for predicting patient disease-specific survival using the c-index. The best performance is highlighted in bold, and the second-best performance is underlined. Over five runs, the standard deviation is reported in brackets. Methods marked * are re-implemented.
@@
PIBD [20] $0.667\pm 0.061$ $0.736\pm 0.072$
PIBD* - MOAB $0.684\pm 0.046$ $\underline{0.749}\pm 0.062$
MMP [22] $0.628\pm 0.064$ $\mathbf{0.753}\pm 0.096$
MMP [22] $0.635\pm 0.064$ $0.738\pm 0.096$
ACMIL* - MOAD-Net [23] $0.658\pm 0.068$ $0.661\pm 0.082$
TransMIL* - MOAD-Net $0.661\pm 0.053$ $0.675\pm 0.068$
SurvPath [13] $0.625\pm 0.056$ $0.655\pm 0.089$
SurvPath* - $p_{ij}$ $0.660\pm 0.047$ $0.665\pm 0.006$
MBFusion [19] --- - $0.644\pm 0.020$
ED-GNN [24] --- - $0.672\pm 0.059$
MoME [36] $\underline{0.686}\pm 0.041$-- -
MuGI [37] $0.681\pm 0.056$-- -
ABMIL MOAD-Net(Ours) $\mathbf{0.691}\pm 0.069$ $0.726\pm 0.049$
Table 4: Survival prediction. Comparison of c-index for BRCA and BLCA datasets using ABMIL and TransMIL models with two fusion stages, late (LF) and dual (DF), across three aggregation methods
BRCA ($\uparrow$)
@@
TransMIL LF $0.606\pm 0.064$ $0.602\pm 0.089$ $0.620\pm 0.115$
DF $0.623\pm 0.032$ $0.645\pm 0.067$ $0.675\pm 0.068$
ABMIL LF $0.625\pm 0.075$ $0.616\pm 0.090$ $0.711\pm 0.095$
DF $0.634\pm 0.023$ $0.640\pm 0.058$ $\mathbf{0.726\pm 0.049}$
BLCA ($\uparrow$)
@@
TransMIL LF $0.599\pm 0.081$ $0.582\pm 0.053$ $0.600\pm 0.108$
DF $0.624\pm 0.068$ $0.595\pm 0.052$ $0.661\pm 0.053$
ABMIL LF $0.558\pm 0.082$ $0.571\pm 0.063$ $0.677\pm 0.054$
DF $0.622\pm 0.063$ $0.561\pm 0.054$ $\mathbf{0.691\pm 0.069}$
Figure 4: Comparison of confusion matrices and t-SNE visualizations for three fusion strategies: (A) MOAD-Net, (B) Late Fusion, and (C) Early Fusion for brain tumor subtyping. The corresponding t-SNE plots are labeled as (A.1), (B.1), and (C.1), respectively.
Figure 5: Visual representation of attention heatmap generated by MOAD-Net for a diffuse glioma, IDH-mutant and 1p19q-retained (astroglial type) tumor (Class 0). (A) The original histology slide is displayed. (B) The heatmap shows areas of high attention (red) and low attention (blue), with regions of diagnostic relevance highlighted. (C) Representative patches with high attention are bordered in red, potentially indicating hallmark features of astroglial differentiation and cellular atypia crucial for diagnosis. (D) Representative patches with low attention are bordered in blue, reflecting regions of reactive gliosis or low tumor infiltration. The color bar illustrates the attention scale from high (red) to low (blue).
5 Discussion
In CNS tumor subtyping, the integration of DNA methylation profiling with WSI data remains relatively underexplored despite its pivotal role in brain tumor diagnosis. This integration holds significant potential to address key challenges, such as differentiating fine-grained CNS tumor subtypes with overlapping morphological features and managing large-scale, heterogeneous datasets. To tackle these issues, our innovative framework, MOAD-Net, incorporates DNA methylation data at multiple stages, enabling a comprehensive fusion approach. To achieve this, we propose using an MLP to map WSI patch features with DNA methylation features. This introduces nonlinearity, allowing the model to learn complex relationships while maintaining a low-dimensional representation, thus avoiding unnecessary complexity. However, we observe that when applying MOAD-Net with complex fusion models like SurvPath and MCAT, performance declines, and training becomes prolonged without effective learning. In contrast, using dual fusion embeddings $\mathbf{p}_{ij}$ as input to these models improves performance without introducing additional complexity. In addition to the extensive ablation studies presented in Table 1, we evaluated MOAD-Net by isolating two components—early fusion and late fusion—using projected feature embeddings, $\mathbf{p}_{ij}$, as input to the ABMIL model. This analysis highlights the role of early fusion in enhancing overall performance and generating interpretable heatmaps. For example, as shown in Fig. 5, the red regions of the heatmap represent a specific subtype (class 0), while the blue patches identify tumor regions that lack the defining characteristics of class 0. This demonstrates the effectiveness of post-fusion embeddings in ABMIL, which successfully separates tumor patches associated with the specific subtype, improving interpretability and performance. MOAD-Net’s use of outer arithmetic operations enables the model to capture diverse relationships between modalities, extending beyond mere additive effects. This enhanced representation is especially beneficial for tasks requiring fine-grained predictions like the complex brain tumor subtyping, as seen in the consistent improvement across all metrics in Table 1. Our proposed approach also appears to scale well across different architectures, as indicated by the strong performance of models such as ACMIL and TransMIL when combined with MOAD-Net.

### v2

@@ 3.2.2 Ablation studies
we conducted three ablation experiments to comprehensively evaluate MOAD-FNet. We first removed the MOAB fusion block, using only ABMIL. Here, $\mathbf{p}_{ij}$ served as the input to ABMIL, and the resulting feature embedding $\mathbf{v}_{i}$ was directly fed to the classifier layer to assess the distinct impact of both ABMIL and the DNA modality. In the second experiment, we removed the early fusion block (illustrated in Fig. 1), making $\mathbf{e}_{ij}$ the input to the gated attention block, which is then followed by MOAD-FNet. Last, we employed a task-agnostic encoder ConvNext.v2 [37], pre-trained on ImageNet, to extract features from the WSI and tested MOAD-FNet with this setup. For the survival prediction task, we evaluated the MOAD-FNet approach using the two most common baseline fusion models: ABMIL [35] and TransMIL [33]. We used the late fusion techniques: concatenation and Kronecker product and compared these against MOAB within both fusion settings.
@@ 3.3 Data and Code Availability
For brain tumor subtyping, we obtained data from NHNN through a rigorous application process to BRAIN UK, securing anonymized H&E slides of tissue samples and epigenetic data. For survival prediction, WSIs are publicly available through the TCGA repository, with corresponding omic data from [13]. Source code will be made available upon acceptance.
@@ 4 Results
Subtyping results. In Tables 1 and 2, we present the results for the subtyping task on the NHNN BRAIN UK dataset. MOAD-FNet integrated with ABMIL consistently demonstrates superior performance in brain tumor subtyping. Specifically, Table 1 shows that the SNN model performs well across metrics using omics-only data, achieving an F1-Macro score of $0.726(\pm 0.003)$, marginally outperforming the MLP. In contrast, WSI-only models show relatively low performance across metrics, with ABMIL achieving an F1-Macro of just $0.247(\pm 0.004)$ and TransMIL performing slightly worse at $0.217(\pm 0.019)$ likely due to the issue of indistinguishable patterns between most subtypes and the presence of underrepresented rare classes. These results indicate that omics data alone offers strong predictive power, however further improvements and interpretability are limited without incorporating WSI data. On the other hand, multimodal models substantially outperformed unimodal ones. For instance, ABMIL-MOAD-FNet achieved the best scores across all metrics, with an F1-Macro of $0.745(\pm 0.025)$ and an F1-Micro of $0.820(\pm 0.031)$. This represents an improvement of $0.027$ in F1-Macro compared to the second-best multimodal model, TransMIL-MOAD-FNet, further emphasizing the effectiveness of the MOAD-FNet architecture in leveraging multimodal data for richer and more informative representations. To assess statistical significance, we performed a Wilcoxon rank-sum test [38] comparing the F1-Macro scores of ABMIL-MOAD-FNet with all other multimodal models in Table 1. The test yielded a $p$-value of $0.043$, indicating a statistically significant difference at the 0.05 significance level. These results demonstrate that integrating MOAD-FNet with advanced multimodal architectures leads to significant performance gains over single-stage fusion methods. This underscores the critical role of dual fusion strategies in effectively combining complementary features from multiple modalities, ultimately driving superior predictive accuracy and robustness.
@@
Table 1: Subtyping prediction Results on the NHNN UK BRAIN dataset. We show MOAD-FNet combined with SOTA baseline models. The best performance is highlighted in bold. Concatenation and Kronecker products are denoted (Cat) and (KP). The gray rows correspond to models integrating MOAD-FNet or using the early fusion embeddings $\mathbf{p}_{ij}$. These demonstrate improved performance compared to standalone baseline models.
Model F1-Macro F1-Micro Precision Recall
@@
ACMIL MOAD-FNet $0.501$ $(0.011)$ $0.729$ $(0.031)$ $0.602$ $(0.032)$ $0.490$ $(0.014)$
TransMIL (Cat) $0.711$ $(0.005)$ $0.799$ $(0.005)$ $0.744$ $(0.003)$ $0.707$ $(0.002)$
TransMIL (KP) $0.483$ $(0.003)$ $0.741$ $(0.016)$ $0.485$ $(0.007)$ $0.506$ $(0.008)$
TransMIL MOAD-FNet $0.724$ $(0.005)$ $0.816$ $(0.002)$ $0.759$ $(0.009)$ $0.719$ $(0.001)$
MCAT $0.402$ $(0.046)$ $0.661$ $(0.021)$ $0.408$ $(0.045)$ $0.414$ $(0.039)$
MCAT $p_{ij}$ $0.432$ $(0.015)$ $0.703$ $(0.017)$ $0.441$ $(0.031)$ $0.446$ $(0.011)$
SURVPATH $0.424$ $(0.008)$ $0.697$ $(0.012)$ $0.478$ $(0.038)$ $0.425$ $(0.013)$
SURVPATH $p_{ij}$ $0.531$ $(0.008)$ $0.761$ $(0.008)$ $0.632$ $(0.005)$ $0.520$ $(0.007)$
ABMIL (Cat) $0.718$ $(0.013)$ $0.806$ $(0.002)$ $0.764$ $(0.018)$ $0.717$ $(0.023)$
ABMIL (KP) $0.447$ $(0.024)$ $0.737$ $(0.020)$ $0.481$ $(0.043)$ $0.464$ $(0.032)$
ABMIL MOAD-FNet 0.745 $(0.025)$ 0.820 $(0.013)$ 0.769 $(0.016)$ 0.745 $(0.035)$
Figure 3: Visual representation of attention heatmap generated by MOAD-FNet for a diffuse glioma, IDH-mutant and 1p19q-retained (astroglial type) tumor (Class 0). (A) The original histology slide is displayed. (B) The heatmap shows areas of high attention (red) and low attention (blue), with regions of diagnostic relevance highlighted. (C) Representative patches with high attention are bordered in red, potentially indicating hallmark features of astroglial differentiation and cellular atypia crucial for diagnosis. (D) Representative patches with low attention are bordered in blue, reflecting regions of low tumor infiltration. The color bar illustrates the attention scale from high (red) to low (blue).
Qualitative results shown in Fig. 5 further illustrate MOAD-FNet’s effectiveness in classifying most tumor subtypes, showcasing the strong impact of its intermingled features. To assess class separation accuracy in the t-SNE representations we calculated silhouette scores for early and late fusion, as they displayed similar patterns. The late fusion t-SNE achieved a silhouette score of 0.33, while MOAD-FNet scored 0.37, indicating that it provides more accurate class separation. This is also evident in the box plot in Fig. 4, where MOAD-FNet displays fewer low outliers and a more compact distribution than early and late fusion.
Figure 4: Comparison of F1-scores across different fusion methods for glioma classification. The radar chart illustrates the F1-score performance across all 20 classes, highlighting distinct patterns for Early Fusion, MOAD-FNet, and Late Fusion. The bar plots zoom in on specific glioblastoma and glioma classes, showing class-level performance variations across fusion methods. The box plot provides a summary of F1-score distributions, showcasing the variability and consistency of each fusion method.
Figure 5: Comparison of confusion matrices and t-SNE visualizations for three fusion strategies: (A) MOAD-FNet, (B) Late Fusion, and (C) Early Fusion for brain tumor subtyping. The corresponding t-SNE plots are labeled as (A.1), (B.1), and (C.1), respectively.
In Fig. 5A, we show the confusion matrix corresponding to our model MOAD-FNet, while in Fig. 5B and C, we see the results for late and early fusion, respectively. These demonstrate that our dual-fusion MOAD-FNet method excels in classifying subtypes, especially for minority classes. Notably, classes 4, 9, and 13, which was misclassified with late fusion (Fig. 5B), were correctly classified by MOAD-FNet. Conversely, late fusion performed better for classes 2 and 16, highlighting the challenges of handling heterogeneous multimodal data. This also underscores the need for tailored fusion techniques to address different data complexities. Furthermore, Fig. 4C1 and C2 illustrate that MOAD-FNet’s dual fusion approach considerably enhances classification of all glioblastoma and diffuse glioma subtypes, even in cases with highly overlapping morphological features.
We present ablation results in Table 2. The findings highlight the necessity for advanced fusion techniques to effectively integrate complementary features across modalities. Notably, the full MOAD-FNet model achieved the highest F1-Macro score of $0.745$ (±$(\pm 0.035)$, demonstrating a notable improvement over both early fusion, with a gain of $0.101$, and late fusion, with a gain of $0.055$. Surprisingly, the task-agnostic ConvNeXt encoder achieved an F1-Macro of0.7320.7320.7320.732 (±$(\pm 0.012)$, marginally lower than the pretrained UNI encoder. This suggests that the core improvement in the full model stems not from the backbone architecture but from the dual fusion mechanism that effectively integrates complementary features across modalities.
@@
Table 2: Subtyping prediction Ablation study of MOAD-FNet showing the performance of early, late, and dual fusion methods using the NHNN BRAIN UK dataset.
Ablation Model Description F1-Macro
ABMIL Early fusion with $\mathbf{p}_{ij}$ $0.644$ $(0.010)$
ABMIL - MOAB Late fusion with $\mathbf{e}_{ij}$ $0.690$ $(0.030)$
MOAD-FNet ConvNeXt encoder $0.732$ $(0.012)$
MOAD-FNet Full model $0.745$ $(0.035)$
Survival prediction results. The results shown in Table 3 demonstrate that integrating MOAD-FNet with existing SOTA methods (indicated by *) derives consistent performance improvement compared to baseline models. For instance, our approach achieves the highest c-index of $0.691(\pm 0.069)$ for BLCA and $0.726(\pm 0.049)$ for BRCA, surpassing all other methods in predicting patient disease-specific survival for BLCA while performing on par with the top-performing model MMP [22] for BRCA. Interestingly, the best results on BRCA were obtained by PIBD with MOAB, achieving a strong c-index of $0.749(\pm 0.062)$ [22].
@@
Table 3: Survival prediction Results of MOAD-FNet with baselines for predicting patient disease-specific survival using the c-index. The best performance is highlighted in bold, and the second-best performance is underlined. Over five runs, the standard deviation is reported in brackets. Methods marked * are re-implemented.
@@
PIBD [20] $0.667$ $(0.061)$ $0.736$ $(0.072)$
PIBD* - MOAB $0.684$ $(0.046)$ $\underline{0.749}$ $(0.062)$
MMP [22] $0.628$ $(0.064)$ $\mathbf{0.753}$ $(0.096)$
MMP [22] $0.635$ $(0.064)$ $0.738$ $(0.096)$
ACMIL* - MOAD-FNet [23] $0.658$ $(0.068)$ $0.661$ $(0.082)$
TransMIL* - MOAD-FNet $0.661$ $(0.053)$ $0.675$ $(0.068)$
SurvPath [13] $0.625$ $(0.056)$ $0.655$ $(0.089)$
SurvPath* - $p_{ij}$ $0.660$ $(0.047)$ $0.665$ $(0.006)$
MBFusion [19] --- - $0.644$ $(0.020)$
ED-GNN [24] --- - $0.672$ $(0.059)$
MoME [39] $\underline{0.686}$ $(0.041)$-- -
MuGI [40] $0.681$ $(0.056)$-- -
ABMIL MOAD-FNet(Ours) $\mathbf{0.691}$ $(0.069)$ $0.726$ $(0.049)$
Note that we did not test MOAD-FNet with MMP [27] because MMP already incorporates two early fusion stages (transformer and optimal transport). Adding MOAB would bring the total to four fusion stages, potentially introducing additional complexity and noise without clear performance benefits.
Table 4: Survival prediction Ablation. Comparison of c-index for TGGA BRCA and BLCA datasets using ABMIL and TransMIL models with two fusion stages, late (LF) and dual (DF), across three aggregation methods.
BRCA c-index ($\uparrow$)
@@
TransMIL LF $0.606$ $(0.064)$ $0.602$ $(0.089)$ $0.620$ $(0.115)$
DF $0.623$ $(0.032)$ $0.645$ $(0.067)$ $0.675$ $(0.068)$
ABMIL LF $0.625$ $(0.075)$ $0.616$ $(0.090)$ $0.711$ $(0.095)$
DF $0.634$ $(0.023)$ $0.640$ $(0.058)$ $\mathbf{0.726}$ $(0.049)$
BLCA c-index ($\uparrow$)
@@
TransMIL LF $0.599$ $(0.081)$ $0.582$ $(0.053)$ $0.600$ $(0.108)$
DF $0.624$ $(0.068)$ $0.595$ $(0.052)$ $0.661$ $(0.053)$
ABMIL LF $0.558$ $(0.082)$ $0.571$ $(0.063)$ $0.677$ $(0.054)$
DF $0.622$ $(0.063)$ $0.561$ $(0.054)$ $\mathbf{0.691}$ $(0.069)$
To further evaluate the effectiveness of MOAD-FNet in the survival prediction task, we conducted additional ablation studies presented in Table 4. The results demonstrate that the ABMIL model with Dual Fusion (DF) consistently delivers superior performance across both the BRCA and BLCA datasets, particularly when paired with the MOAB aggregation method. For the BRCA dataset, ABMIL-DF with MOAB achieves the highest c-index of $0.726(\pm 0.049)$, outperforming TransMIL-DF with MOAB. Similarly, for the BLCA dataset, ABMIL-DF with MOAB achieves the highest c-index of $0.691(\pm 0.069)$, surpassing TransMIL-DF with MOAB. The Late Fusion (LF) results follow a similar pattern, with ABMIL consistently outperforming TransMIL across all aggregation methods. These findings underscore the effectiveness of ABMIL, particularly with the Dual Fusion strategy and MOAB aggregation.
Limitations. While our method demonstrates promising performance across multiple tasks, it’s important to acknowledge some limitations. MOAD-FNet faces two primary challenges: first, balancing early and late fusion presents a trade-off; although MOAD-FNet leverages both stages to capture the strengths of each, finding the optimal balance can be complex, with some cases exhibiting performance variations depending on the fusion stage. Second, our late fusion block incorporating MOAB operates on latent space vectors derived from the early fusion stage, where omics features have already been blended. Consequently, identifying the specific CpG feature with the most profound impact on the outcome becomes challenging.